Tal is making a nut mixture to sell at the local farmers' market. He mixes 5 pounds of pistachios with a nut mixture that is 70% pistachios. The resulting mixture is 80% pistachios. How many pounds of nut mixture does Tal make?
Answer by ptaylor(2198)

Let x=number of lbs of nut mixture used
Now we know that the amount of pistachios that we have before the mixture takes place has to equal the amount of pistachios that we have after the mixture takes place.
Amount of Pistachios before the mixture takes place: 5+0.70x
Amount of pistachios after the mixture takes place: 0.80(5+x)
So, our equation to solve is:
5+0.70x=0.80(5+x) get rid of parens and simplify
5+0.70x=4+0.80x subtract 5 and also 0.80x from each side
-0.10x=-1
x=10 lbs------amount of nut mixture needed
CK
5+0.70*10=0.80*15
5+7=12
12=12
Hope this helps---ptaylor